# Motives of nullcones of quiver representations

### v2

@@ 3. Applications and examples
We have the following equivalent form of the recursion:
$[N_{\bf d}]=-\sum_{{\bf e}\lneqq{\bf d}}(-1)^{\dim{\bf d}-\dim{\bf e}}\cdot\mathbb{L}^{\sum_{i}({d_{i}\choose 2}-{e_{i}\choose 2})-\langle{\bf e},{\bf d}-{\bf e}\rangle}\cdot\left[{{\bf d}\atop{\bf e}}\right]_{\mathbb{L}}\cdot{[N_{\bf e}]}.$
In particular, there exist polynomials $n_{\bf d}(x)\in\mathbb{Z}[x]$ such that $[N_{\bf d}]=n_{\bf d}(\mathbb{L})$.
@@
The recursion follows from Theorem 2.2 just by making the motive of $G_{\bf e}$ explicit. Defining polynomials $n_{\bf d}$ by the same recursion with $\mathbb{L}$ replaced by $x$, and starting with $n_{0}=1$, we obviously have $n_{\bf d}\in\mathbb{Z}[x^{\pm 1}]$, and $n_{\bf d}(\mathbb{L})=[N_{\bf d}]$. By the definition of the Euler form, the $\mathbb{L}$-exponent in the recursion is always nonnegative, thus we have $n_{\bf d}(x)\in\mathbb{Z}[x]$.∎
We define the ring $R_{\mathbb{L}}\llbracket Q_{0}\rrbracket$ as the twisted formal power series ring with topological $R$-basis elements $t^{\bf d}$ for ${\bf d}\in\mathbb{N}Q_{0}$ and multiplication
$t^{\bf d}\cdot t^{\bf e}=\mathbb{L}^{-\langle{\bf d},{\bf e}\rangle}\cdot t^{{\bf d}+{\bf e}}.$
@@
In $R_{\mathbb{L}}\llbracket Q_{0}\rrbracket$, we have the identity
@@
Expanding the left hand side of the claimed equation, its $t^{\bf d}$-coefficient equals
@@
which equals 00 for ${\bf d}\not=0$ by Theorem 2.2.∎
Using this identity of generating functions, we can resolve the recursion of Theorem 2.2:
@@
$[N_{\bf d}]=(-1)^{\dim{\bf d}}\cdot\mathbb{L}^{\sum_{i}{d_{i}\choose 2}}\cdot\sum_{{\bf d}^{*}}(-1)^{s}\cdot\mathbb{L}^{-\sum_{k<l}\langle{\bf d}^{k},{\bf d}^{l}\rangle}\cdot\left[{{\bf d}\atop{{\bf d}^{1},\ldots,{\bf d}^{s}}}\right]_{\mathbb{L}},$
where the sum ranges over all ordered decompositions ${\bf d}={\bf d}^{1}+\ldots+{\bf d}^{s}$ of ${\bf d}$ into non-zero parts ${\bf d}^{k}$.
@@
$\sum_{\bf d}\frac{[N_{\bf d}]}{[G_{\bf d}]}t^{\bf d}=(1+\sum_{\bf d\not={0}}\frac{t^{\bf d}}{(\mathbb{L})_{\bf d}})^{-1}=\sum_{s\geq 0}(-1)^{s}(\sum_{{\bf d}\not=0}\frac{t^{\bf d}}{(\mathbb{L})_{\bf d}})^{s}=$
=$\sum_{s\geq 0}(-1)^{s}\cdot\sum_{{\bf d}^{1},\ldots,{\bf d}^{s}\not=0}\mathbb{L}^{-\sum_{k<l}\langle{\bf d}^{k},{\bf d}^{l}\rangle}\frac{t^{\bf d}}{(\mathbb{L})_{{\bf d}^{1}}\cdot\ldots\cdot(\mathbb{L})_{{\bf d}^{s}}}.$
Comparing coefficients of $t^{\bf d}$ on both sides and rewriting $[G_{\bf d}]$ using the Pochhammer symbol, we arrive at the claimed formula.∎
We call a quiver symmetric if its Euler form is symmetric. Equivalently, denoting by $r_{ij}$ the number of arrows from $i$ to $j$ in $Q$ for $i,j\in Q_{0}$, symmetry is given by $r_{ij}=r_{ji}$ for all $i,j\in Q_{0}$. In this case, it is immediately verified that
$x^{\bf d}=(-\mathbb{L}^{1/2})^{-\langle{\bf d},{\bf d}\rangle}t^{\bf d}$
fulfills $x^{\bf d}\cdot x^{\bf e}=x^{{\bf d}+{\bf e}}$; consequently, we do not have to work with twisted multiplications of formal series in the symmetric case. We extend the base ring of motives by adjoinging a square root of $\mathbb{L}$. This allows us to define motivic Donaldson–Thomas invariants (see [14] for an exposition) by the identity
$\sum_{\mathbf{d}}(-\mathbb{L}^{1/2})^{\langle\mathbf{d},\mathbf{d}\rangle}\frac{[R_{\mathbf{d}}]}{[G_{\mathbf{d}}]}x^{\mathbf{d}}={\rm Exp}(\frac{1}{\mathbb{L}^{-1/2}-\mathbb{L}^{1/2}}\sum_{\mathbf{d}\not=0}{\rm DT}^{Q}_{\mathbf{d}}(\mathbb{L})x^{\mathbf{d}})$
in $R[\mathbb{L}^{1/2}]\llbracket Q_{0}\rrbracket$. Here ${\rm Exp}$ is defined by
${\rm Exp}(\sum_{{\bf d}\not=0}\sum_{k}c_{{\bf d},k}q^{k/2}x^{\bf d})=\prod_{{\bf d}\not=0}\prod_{k}(1-q^{k/2}x^{\bf d})^{-c_{{\bf d},k}}.$
It is known by [3] that the Donaldson–Thomas invariant is a polynomial in $-\mathbb{L}^{1/2}$ with nonnegative coefficients, that is, ${\rm DT}^{Q}_{\mathbf{d}}(\mathbb{L})\in\mathbb{Z}[\mathbb{L}^{\pm 1/2}]$.
@@
$\sum_{\mathbf{d}}(-\mathbb{L}^{1/2})^{\langle\mathbf{d},\mathbf{d}\rangle}\frac{[N_{\mathbf{d}}]}{[G_{\mathbf{d}}]}x^{\mathbf{d}}={\rm Exp}(\frac{1}{\mathbb{L}^{-1/2}-\mathbb{L}^{1/2}}\sum_{\mathbf{d}\not=0}{\rm DT}^{Q}_{\bf d}(\mathbb{L}^{-1})x^{\mathbf{d}}).$
@@
The left hand side of the identity defining the DT invariants equals
$\sum_{\bf d}(-\mathbb{L}^{1/2})^{-\langle{\bf d},{\bf d}\rangle}\frac{x^{\bf d}}{(\mathbb{L}^{-1})_{\bf d}}.$
Polynomiality of the Donaldson-Thomas invariants allows us to formally interchange $\mathbb{L}$ and $\mathbb{L}^{-1}$, and thus we have
$\sum_{\bf d}(-\mathbb{L}^{1/2})^{\langle{\bf d},{\bf d}\rangle}\frac{x^{\bf d}}{(\mathbb{L})_{\bf d}}={\rm Exp}(\frac{1}{\mathbb{L}^{1/2}-\mathbb{L}^{-1/2}}\sum_{{\bf d}\not=0}{\rm DT}^{Q}_{\bf d}(\mathbb{L}^{-1})x^{\bf d}).$
@@
$\sum_{\mathbf{d}}(-\mathbb{L}^{1/2})^{\langle\mathbf{d},\mathbf{d}\rangle}\frac{[N_{\mathbf{d}}]}{[G_{\mathbf{d}}]}x^{\mathbf{d}}=(\sum_{\bf d}(-\mathbb{L}^{1/2})^{\langle{\bf d},{\bf d}\rangle}\frac{x^{\bf d}}{(\mathbb{L})_{\bf d}})^{-1}.$
@@
$n_{\bf d}(x)=(-1)^{\dim{\bf d}}\cdot{x}^{\sum_{i}{d_{i}\choose 2}}\cdot\sum_{{\bf d}^{*}}(-1)^{s}\cdot{x}^{-\sum_{k<l}\langle{\bf d}^{k},{\bf d}^{l}\rangle}\cdot\left[{{\bf d}\atop{{\bf d}^{1},\ldots,{\bf d}^{s}}}\right]_{x}.$
@@
$-\sum_{k<l}\langle{\bf d}^{k},{\bf d}^{l}\rangle+\sum_{k<l}\sum_{i}d^{k}_{i}d^{l}_{i}=\sum_{k<l}\sum_{i,j}r_{ij}d^{k}_{i}d^{l}_{j}=\frac{1}{2}(\sum_{i,j}r_{ij}d_{i}d_{j}-\sum_{i,j}\sum_{k}r_{ij}d^{k}_{i}d^{k}_{j}).$
Refining an ordered partition ${\bf d}^{*}$ by replacing a part ${\bf d}^{k}$ by two parts ${\bf d}^{\prime}$, ${\bf d}^{\prime\prime}$ (in this order) thus increases the degree of the corresponding summand by $\sum_{i,j}r_{ij}d^{\prime}_{i}d^{\prime\prime}_{j}$. The degree can thus be maximal only for maximally refined decompositions, which are of the form ${\bf i}_{1},\ldots,{\bf i}_{s}$, where each ${\bf i}$ for $i\in Q_{0}$ appears $d_{i}$-times in the sequence; in particular, $s=\dim{\bf d}$. We define
$K_{i}=\{k=1,\ldots,s\,|\,{\bf i}_{k}=i\}$
for $i\in Q_{0}$. The summand corresponding to the sequence ${\bf i}_{1},\ldots,{\bf i}_{s}$ is then of degree
$\frac{1}{2}(\sum_{i,j}r_{ij}d_{i}d_{j}-\sum_{i,j}\sum_{k}r_{ij}\delta_{k\in K_{i}}\delta_{k\in K_{j}})=$
$=\frac{1}{2}(\sum_{i,j}r_{ij}d_{i}d_{j}-\sum_{i}\sum_{k}r_{ii})=\sum_{i}r_{ii}{d_{i}\choose 2}+\sum_{i<j}r_{ij}d_{i}d_{j}.$
All of the corresponding summands have the same sign $(-1)^{\dim{\bf d}}$, thus there are no cancellations in the summation of these top degree terms. The number of these summands equals the number of decompositions of $\{1,\ldots,\dim{\bf d}\}$ into parts $K_{i}$ for $i\in Q_{0}$, which is the quotient of factorials of the claimed formula. Adding the contribution $\sum_{i}{d_{i}\choose 2}$ of the prefactor in the above formula for $n_{\bf d}(x)$ to the degree, we arrive at the claimed formula for the leading term.∎
Remark: A similar proof appears in [13, Theorem 4.1]. Refining the argument there shows that, in case $Q$ is symmetric, the variety $N_{\bf d}$ is equidimensional of dimension
@@
Examples 3.6.
We finish this section with several examples illustrating the previous formulas:
@@
    If $Q$ is acyclic, every representation is nilpotent, thus
    $[N_{\bf d}]=[R_{\bf d}]=\mathbb{L}^{\sum_{\alpha:i\rightarrow j}d_{i}d_{j}}$
    trivially, but this is not obvious from any of the previous formulas. We will use two variants of the $q$-binomial theorem to derive this fact from Corollary 3.2. Namely, we have
    $\sum_{d\geq 0}\frac{t^{d}}{(q)_{d}}=\prod_{k\geq 0}(1-q^{k}t)^{-1},\;\sum_{d\geq 0}\frac{q^{d\choose 2}t^{d}}{(q)_{d}}=\prod_{k\geq 0}(1+q^{k}t),$
@@
    $(\sum_{d\geq 0}\frac{q^{d\choose 2}t^{d}}{(q)_{d}})^{-1}=\prod_{k\geq 0}(1+q^{k}t)^{-1}=\sum_{d\geq 0}\frac{(-t)^{d}}{(q)_{d}}.$
    Since $Q$ is acyclic, we can enumerate its vertices $Q_{0}=\{i_{1},\ldots,i_{n}\}$ such that the existence of an arrow $i_{k}\rightarrow i_{l}$ implies $k<l$. By the definition of the twisted multiplication in $R_{\mathbb{L}}\llbracket Q_{0}\rrbracket$, we then have
    $t_{i_{1}}^{d_{i_{1}}}\cdot\ldots\cdot t_{i_{n}}^{d_{i_{n}}}=\mathbb{L}^{-\langle{\bf d},{\bf d}\rangle+\sum_{k}d_{i_{k}}(d_{i_{k}}+1)/2}t^{\bf d},$
    $t_{i_{n}}^{d_{i_{n}}}\cdot\ldots\cdot t_{i_{1}}^{d_{i_{1}}}=\mathbb{L}^{-\sum_{k}d_{i_{k}}(d_{i_{k}}-1)/2}t^{\bf d}.$
@@
    $(\sum_{\bf d}\frac{t^{\bf d}}{(\mathbb{L})_{\bf d}})^{-1}=\prod_{k=1}^{n}\sum_{d\geq 0}\frac{(-t_{i_{k}})^{d}}{(\mathbb{L})_{d}}=$
    $=\sum_{\bf d}(-1)^{\dim{\bf d}}\frac{\mathbb{L}^{-\langle{\bf d},{\bf d}\rangle+\sum_{i}d_{i}(d_{i}+1)/2}}{(\mathbb{L})_{\bf d}}t^{\bf d}=\sum_{\bf d}\frac{[R_{\bf d}]}{[G_{\bf d}]}t^{\bf d}.$
    Now $[N_{\bf d}]=[R_{\bf d}]$ follows from Corollary 3.2.
@@
    Similarly, if $Q$ is the quiver with a single vertex and one loop, Corollary 3.2, together with the above two versions of the $q$-binomial theorem, easily yields
    $[N_{d}]=\mathbb{L}^{d(d-1)}$
    for $N_{d}$ the variety of nilpotent $d\times d$-matrices. This provides a motivic analogue of the classical result [4].
@@
    $[N_{\bf 1}]=(-1)^{Q_{0}}\cdot\sum_{I_{*}}(-1)^{s}\cdot\mathbb{L}^{\sum_{k<l}|I_{k}\rightarrow I_{l}|},$
    where the sum ranges over all ordered disjoint unions $Q_{0}=\bigcup_{k=1}^{s}I_{k}$, and $|I_{k}\rightarrow I_{l}|$ denotes the number of arrows from vertices in $I_{k}$ to vertices in $I_{l}$. On the other hand, denote by $Q_{1}$ the set of arrows in $Q$, and define the support of a representation $V\in R_{\bf 1}$ as the subset of $Q_{1}$ of all arrows $\alpha$ such that $V_{\alpha}\not=0$. Then $V$ is nilpotent if and only if its support is an acyclic subset $F\subset Q_{1}$, that is, such that the corresponding subquiver is acyclic. For any such acyclic subset, the locally closed subset of $R_{\bf 1}$ of representations having support $F$ is isomorphic to a torus of rank $|F|$, thus has motive $(\mathbb{L}-1)^{|F|}$. We thus arrive at the formula
    $[N_{\bf 1}]=\sum_{{F\subset Q_{1}}\atop\mbox{\tiny acyclic}}(\mathbb{L}-1)^{|F|}.$
@@
    For the quiver $Q$ with a single vertex and $m$ loops, the variety $N_{\bf d}$ consists of tuples of $d\times d$-matrices $(A_{1},\ldots,A_{m})$ such that every polynomial in these matrices is nilpotent, or, equivalently, which can be simultaneously conjugated to strictly upper triangular matrices. We list $[N_{d}]$ for small $d$ using Corollary 3.3:
    $[N_{1}]=1,\;\;\;[N_{2}]=\mathbb{L}^{m+1}+\mathbb{L}^{m}-\mathbb{L},$
    $[N_{3}]=\mathbb{L}^{3m+3}+2\mathbb{L}^{3m+2}+2\mathbb{L}^{3m+1}+\mathbb{L}^{3m}-$
    $-2\mathbb{L}^{2m+3}-2\mathbb{L}^{2m+2}-2\mathbb{L}^{2m+1}+\mathbb{L}^{3},$
    $[N_{4}]=\mathbb{L}^{6m+6}+3\mathbb{L}^{6m+5}+5\mathbb{L}^{6m+4}+6\mathbb{L}^{6m+3}+5\mathbb{L}^{6m+2}+3\mathbb{L}^{6m+1}+\mathbb{L}^{6m}-$
    $-3\mathbb{L}^{5m+6}-6\mathbb{L}^{5m+5}-9\mathbb{L}^{5m+4}--9\mathbb{L}^{5m+3}-6\mathbb{L}^{5m+2}-3\mathbb{L}^{5m+1}+$
    $+\mathbb{L}^{4m+6}+\mathbb{L}^{4m+5}+2\mathbb{L}^{4m+4}+\mathbb{L}^{4m+3}+\mathbb{L}^{4m+2}+$
    $+2\mathbb{L}^{3m+6}+2\mathbb{L}^{3m+5}+2\mathbb{L}^{3m+4}+2\mathbb{L}^{3m+3}-\mathbb{L}^{6}.$

### v3

@@ 3.1 Applications
We have the following equivalent form of the recursion, which holds as an identity in $K_{0}({\rm Var}_{\mathbb{C}})$:
$N_{\bf d}]=-\sum_{{\bf e}\lneqq{\bf d}}(-1)^{\dim{\bf d}-\dim{\bf e}}\cdot\mathbb{L}^{\sum_{i}\big({d_{i}\choose 2}-{e_{i}\choose 2}\big)-\langle{\bf e},{\bf d}-{\bf e}\rangle}\cdot\left[{{\bf d}\atop{\bf e}}\right]_{\mathbb{L}}\cdot{[N_{\bf e}]}.$
In particular, there exist polynomials $n_{\bf d}(x)\in\mathbb{Z}[x]$ such that $N_{\bf d}]=n_{\bf d}(\mathbb{L})$.
@@
The recursion follows from Theorem 2.2 just by making the motive of $G_{\bf e}$ explicit. Defining polynomials $n_{\bf d}$ by the same recursion with $\mathbb{L}$ replaced by $x$, and starting with $n_{0}=1$, we obviously have $n_{\bf d}\in\mathbb{Z}\big[x^{\pm 1}\big]$, and $n_{\bf d}(\mathbb{L})=[N_{\bf d}]$. By the definition of the Euler form, the $\mathbb{L}$-exponent in the recursion is always nonnegative, thus we have $n_{\bf d}(x)\in\mathbb{Z}[x]$.∎
We define the ring $\mathbb{M}_{\mathbb{L}}\llbracket Q_{0}\rrbracket$ as the twisted formal power series ring with topological $\mathbb{M}$-basis elements $t^{\bf d}$ for ${\bf d}\in\mathbb{N}Q_{0}$ and multiplication $t^{\bf d}\cdot t^{\bf e}=\mathbb{L}^{-\langle{\bf d},{\bf e}\rangle}\cdot t^{{\bf d}+{\bf e}}$. Such twists are frequently used for identities of generating series of motives, see, for example, [11, Section 2.1] and [12, Definition 3.2].
@@
In $\mathbb{M}_{\mathbb{L}}\llbracket Q_{0}\rrbracket$, we have the identity
@@
Expanding the left-hand side of the claimed equation, its $t^{\bf d}$-coefficient equals
@@
which equals 0 for ${\bf d}\not=0$ by Theorem 2.2.∎
Using this identity of generating functions, we can resolve the recursion of Theorem 2.2.
@@
$N_{\bf d}]=(-1)^{\dim{\bf d}}\cdot\mathbb{L}^{\sum_{i}{d_{i}\choose 2}}\cdot\sum_{{\bf d}^{*}}(-1)^{s}\cdot\mathbb{L}^{-\sum_{k<l}\langle{\bf d}^{k},{\bf d}^{l}\rangle}\cdot\left[{{\bf d}\atop{{\bf d}^{1},\dots,{\bf d}^{s}}}\right]_{\mathbb{L}},$
where the sum ranges over all ordered decompositions ${\bf d}={\bf d}^{1}+\dots+{\bf d}^{s}$ of $\bf d$} into non-zero parts ${\bf d}^{k}$.
@@
$\sum_{\bf d}\frac{[N_{\bf d}]}{[G_{\bf d}]}t^{\bf d}$ $=\biggl(1+\sum_{\bf d\not={0}}\frac{t^{\bf d}}{(\mathbb{L})_{\bf d}}\biggr)^{-1}=\sum_{s\geq 0}(-1)^{s}\biggl(\sum_{{\bf d}\not=0}\frac{t^{\bf d}}{(\mathbb{L})_{\bf d}}\biggr)^{s}$
$=\sum_{s\geq 0}(-1)^{s}\cdot\sum_{{\bf d}^{1},\dots,{\bf d}^{s}\not=0}\mathbb{L}^{-\sum_{k<l}\langle{\bf d}^{k},{\bf d}^{l}\rangle}\frac{t^{\bf d}}{(\mathbb{L})_{{\bf d}^{1}}\cdots(\mathbb{L})_{{\bf d}^{s}}}.$
Comparing coefficients of $t^{\bf d}$ on both sides and rewriting $G_{\bf d}]$ using the Pochhammer symbol, we arrive at the claimed formula.∎
We call a quiver symmetric if its Euler form is symmetric. Equivalently, denoting by $r_{ij}$ the number of arrows from $i$ to $j$ in $Q$ for $i,j\in Q_{0}$, symmetry is given by $r_{ij}=r_{ji}$ for all $i,j\in Q_{0}$}. In this case, it is immediately verified that $x^{\bf d}=\bigl(-\mathbb{L}^{1/2}\bigr)^{-\langle{\bf d},{\bf d}\rangle}t^{\bf d}$ fulfills $x^{\bf d}\cdot x^{\bf e}=x^{{\bf d}+{\bf e}}$; consequently, we do not have to work with twisted multiplications of formal series in the symmetric case. We extend the base ring of motives by adjoinging a square root of $\mathbb{L}$. This allows us to define motivic Donaldson–Thomas invariants (see [14] for an exposition) by the identity
$\sum_{\mathbf{d}}\bigl(-\mathbb{L}^{1/2}\bigr)^{\langle\mathbf{d},\mathbf{d}\rangle}\frac{[R_{\mathbf{d}}]}{[G_{\mathbf{d}}]}x^{\mathbf{d}}=\operatorname{Exp}\biggl(\frac{1}{\mathbb{L}^{-1/2}-\mathbb{L}^{1/2}}\sum_{\mathbf{d}\not=0}{\rm DT}^{Q}_{\mathbf{d}}(\mathbb{L})x^{\mathbf{d}}\biggr)$
in $\mathbb{M}\big[\mathbb{L}^{1/2}\big]\llbracket Q_{0}\rrbracket$. Here $\operatorname{Exp}$ is defined by
$\operatorname{Exp}\biggl(\sum_{{\bf d}\not=0}\sum_{k}c_{{\bf d},k}q^{k/2}x^{\bf d}\biggr)=\prod_{{\bf d}\not=0}\prod_{k}\bigl(1-q^{k/2}x^{\bf d}\bigr)^{-c_{{\bf d},k}}.$
It is known by [3] that the Donaldson–Thomas invariant is a polynomial in $\mathbb{L}^{1/2}$ with nonnegative coefficients, that is, ${\rm DT}^{Q}_{\mathbf{d}}(\mathbb{L})\in\mathbb{Z}\big[\mathbb{L}^{\pm 1/2}\big]$.
@@
$\sum_{\mathbf{d}}\bigl(-\mathbb{L}^{1/2}\bigr)^{\langle\mathbf{d},\mathbf{d}\rangle}\frac{[N_{\mathbf{d}}]}{[G_{\mathbf{d}}]}x^{\mathbf{d}}=\operatorname{Exp}\biggl(\frac{1}{\mathbb{L}^{-1/2}-\mathbb{L}^{1/2}}\sum_{\mathbf{d}\not=0}{\rm DT}^{Q}_{\bf d}\bigl(\mathbb{L}^{-1}\bigr)x^{\mathbf{d}}\biggr).$
@@
The left-hand side of the identity defining the DT invariants equals
$\sum_{\bf d}\bigl(-\mathbb{L}^{1/2}\bigr)^{-\langle{\bf d},{\bf d}\rangle}\frac{x^{\bf d}}{\bigl(\mathbb{L}^{-1}\bigr)_{\bf d}}.$
Polynomiality of the Donaldson–Thomas invariants allows us to formally interchange $\mathbb{L}$ and $\mathbb{L}^{-1}$, and thus we have
$\sum_{\bf d}\bigl(-\mathbb{L}^{1/2}\bigr)^{\langle{\bf d},{\bf d}\rangle}\frac{x^{\bf d}}{(\mathbb{L})_{\bf d}}=\operatorname{Exp}\biggl(\frac{1}{\mathbb{L}^{1/2}-\mathbb{L}^{-1/2}}\sum_{{\bf d}\not=0}{\rm DT}^{Q}_{\bf d}\bigl(\mathbb{L}^{-1}\bigr)x^{\bf d}\biggr).$
@@
$\sum_{\mathbf{d}}\bigl(-\mathbb{L}^{1/2}\bigr)^{\langle\mathbf{d},\mathbf{d}\rangle}\frac{[N_{\mathbf{d}}]}{[G_{\mathbf{d}}]}x^{\mathbf{d}}=\biggl(\sum_{\bf d}\bigl(-\mathbb{L}^{1/2}\bigr)^{\langle{\bf d},{\bf d}\rangle}\frac{x^{\bf d}}{(\mathbb{L})_{\bf d}}\biggr)^{-1}.$
@@
$n_{\bf d}(x)=(-1)^{\dim{\bf d}}\cdot{x}^{\sum_{i}{d_{i}\choose 2}}\cdot\sum_{{\bf d}^{*}}(-1)^{s}\cdot{x}^{-\sum_{k<l}\langle{\bf d}^{k},{\bf d}^{l}\rangle}\cdot\left[{{\bf d}\atop{{\bf d}^{1},\dots,{\bf d}^{s}}}\right]_{x}.$
@@
$\sum_{k<l}\big\langle{\bf d}^{k},{\bf d}^{l}\big\rangle+\sum_{k<l}\sum_{i}d^{k}_{i}d^{l}_{i}=\sum_{k<l}\sum_{i,j}r_{ij}d^{k}_{i}d^{l}_{j}=\frac{1}{2}\biggl(\sum_{i,j}r_{ij}d_{i}d_{j}-\sum_{i,j}\sum_{k}r_{ij}d^{k}_{i}d^{k}_{j}\biggr).$
Refining an ordered partition $\bf d$$^{*}$ by replacing a part ${\bf d}^{k}$ by two parts $\bf d$$^{\prime}$, $\bf d$$^{\prime\prime}$ (in this order) thus increases the degree of the corresponding summand by $\sum_{i,j}r_{ij}d^{\prime}_{i}d^{\prime\prime}_{j}$. The degree can thus be maximal only for maximally refined decompositions, which are of the form ${\bf i}_{1},\dots,{\bf i}_{s}$, where each $\bf i$} for $i\in Q_{0}$ appears $d_{i}$-times in the sequence; in particular, $s=\dim{\bf d}$. We define
$K_{i}=\{k=1,\dots,s\mid{\bf i}_{k}=i\}$
for $i\in Q_{0}$. The summand corresponding to the sequence ${\bf i}_{1},\dots,{\bf i}_{s}$ is then of degree
$\frac{1}{2}\biggl(\sum_{i,j}r_{ij}d_{i}d_{j}-\sum_{i,j}\sum_{k}r_{ij}\delta_{k\in K_{i}}\delta_{k\in K_{j}}\biggr)$ $=\frac{1}{2}\biggl(\sum_{i,j}r_{ij}d_{i}d_{j}-\sum_{i}\sum_{k}r_{ii}\biggr)$
$=\sum_{i}r_{ii}{d_{i}\choose 2}+\sum_{i<j}r_{ij}d_{i}d_{j}.$
All of the corresponding summands have the same sign $(-1)^{\dim{\bf d}}$, thus there are no cancellations in the summation of these top degree terms. The number of these summands equals the number of decompositions of $\{1,\dots,\dim{\bf d}\}$ into parts $K_{i}$ for $i\in Q_{0}$, which is the quotient of factorials of the claimed formula. Adding the contribution $\sum_{i}{d_{i}\choose 2}$ of the prefactor in the above formula for $n_{\bf d}(x)$ to the degree, we arrive at the claimed formula for the leading term.∎
Remark 3.6.
A similar proof appears in [13, Theorem 4.1]. Refining the argument there shows that, in case $Q$ is symmetric, the variety $N_{\bf d}$ is equidimensional of dimension
@@
3.2 Examples
We finish this section with several examples illustrating the previous formulas.
@@ 3.2 Examples
    If $Q$ is acyclic, every representation is nilpotent, thus $N_{\bf d}]=[R_{\bf d}]=\mathbb{L}^{\sum_{\alpha\colon i\rightarrow j}d_{i}d_{j}}$ trivially, but this is not obvious from any of the previous formulas. We will use two variants of the $q$-binomial theorem to derive this fact from Corollary 3.2. Namely, we have
    $\sum_{d\geq 0}\frac{t^{d}}{(q)_{d}}=\prod_{k\geq 0}\bigl(1-q^{k}t\bigr)^{-1},\sum_{d\geq 0}\frac{q^{d\choose 2}t^{d}}{(q)_{d}}=\prod_{k\geq 0}\bigl(1+q^{k}t\bigr),$
@@
    $\left(\sum_{d\geq 0}\frac{q^{d\choose 2}t^{d}}{(q)_{d}}\right)^{-1}=\prod_{k\geq 0}\bigl(1+q^{k}t\bigr)^{-1}=\sum_{d\geq 0}\frac{(-t)^{d}}{(q)_{d}}.$
    Since $Q$ is acyclic, we can enumerate its vertices $Q_{0}=\{i_{1},\dots,i_{n}\}$ such that the existence of an arrow $i_{k}\rightarrow i_{l}$ implies $k<l$. By the definition of the twisted multiplication in $\mathbb{M}_{\mathbb{L}}\llbracket Q_{0}\rrbracket$, we then have
    $t_{i_{1}}^{d_{i_{1}}}\cdots t_{i_{n}}^{d_{i_{n}}}=\mathbb{L}^{-\langle{\bf d},{\bf d}\rangle+\sum_{k}d_{i_{k}}(d_{i_{k}}+1)/2}t^{\bf d},\qquad t_{i_{n}}^{d_{i_{n}}}\cdots t_{i_{1}}^{d_{i_{1}}}=\mathbb{L}^{-\sum_{k}d_{i_{k}}(d_{i_{k}}-1)/2}t^{\bf d}.$
@@
    $\left(\sum_{\bf d}\frac{t^{\bf d}}{(\mathbb{L})_{\bf d}}\right)^{-1}$ $=\prod_{k=1}^{n}\sum_{d\geq 0}\frac{(-t_{i_{k}})^{d}}{(\mathbb{L})_{d}}=\sum_{\bf d}(-1)^{\dim{\bf d}}\frac{\mathbb{L}^{-\langle{\bf d},{\bf d}\rangle+\sum_{i}d_{i}(d_{i}+1)/2}}{(\mathbb{L})_{\bf d}}t^{\bf d}$
    $=\sum_{\bf d}\frac{[R_{\bf d}]}{[G_{\bf d}]}t^{\bf d}.$
    Now $N_{\bf d}]=[R_{\bf d}]$ follows from Corollary 3.2.
@@
    Similarly, if $Q$ is the quiver with a single vertex and one loop, Corollary 3.2, together with the above two versions of the $q$-binomial theorem, easily yields $N_{d}]=\mathbb{L}^{d(d-1)}$ for $N_{d}$ the variety of nilpotent $d\times d$ matrices. This provides a motivic analogue of the classical result [4].
@@
    $N_{\bf 1}]=(-1)^{Q_{0}}\cdot\sum_{I_{*}}(-1)^{s}\cdot\mathbb{L}^{\sum_{k<l}|I_{k}\rightarrow I_{l}|},$
    where the sum ranges over all tuples $I_{*}=(I_{1},\dots,I_{s})$ of subsets of $Q_{0}$ inducing a disjoint union $Q_{0}=\bigcup_{k=1}^{s}I_{k}$, and $|I_{k}\rightarrow I_{l}|$ denotes the number of arrows from vertices in $I_{k}$ to vertices in $I_{l}$. On the other hand, denote by $Q_{1}$ the set of arrows in $Q$, and define the support of a representation $V\in R_{\bf 1}$ as the subset of $Q_{1}$ of all arrows $\alpha$ such that $V_{\alpha}\not=0$. Then $V$ is nilpotent if and only if its support is an acyclic subset $F\subset Q_{1}$, that is, such that the corresponding subquiver is acyclic. For any such acyclic subset, the locally closed subset of $R_{\bf 1}$ of representations having support $F$ is isomorphic to a torus of rank $|F|$, thus has motive $(\mathbb{L}-1)^{|F|}$ (as an aside, we note that every irreducible component of $N_{\bf 1}$ is thus toric). We then arrive at the formula $N_{\bf 1}]=\sum_{{F\subset Q_{1}}\atop\mbox{\tiny acyclic}}(\mathbb{L}-1)^{|F|}$.
@@
    For the quiver $Q$ with a single vertex and $m$ loops, the variety $N_{\bf d}$ consists of tuples of $d\times d$ matrices $(A_{1},\dots,A_{m})$ such that every polynomial in these matrices is nilpotent, or, equivalently, which can be simultaneously conjugated to strictly upper triangular matrices. We list $N_{d}]$ for small $d$ using Corollary 3.3
    $[N_{1}]=1,\qquad[N_{2}]=\mathbb{L}^{m+1}+\mathbb{L}^{m}-\mathbb{L},$
    $[N_{3}]=\mathbb{L}^{3m+3}+2\mathbb{L}^{3m+2}+2\mathbb{L}^{3m+1}+\mathbb{L}^{3m}-2\mathbb{L}^{2m+3}-2\mathbb{L}^{2m+2}-2\mathbb{L}^{2m+1}+\mathbb{L}^{3},$
    $[N_{4}]=\mathbb{L}^{6m+6}+3\mathbb{L}^{6m+5}+5\mathbb{L}^{6m+4}+6\mathbb{L}^{6m+3}+5\mathbb{L}^{6m+2}+3\mathbb{L}^{6m+1}+\mathbb{L}^{6m}$
[…]